For a new lab test you need 75 ml of a 2.25 M solution of CuSO4. How much CuSO4 do you need?
Answer #1

volume , V = 75 mL
= 7.5*10^-2 L


use:
number of mol,
n = Molarity * Volume
= 2.25*7.5*10^-2
= 0.1687 mol

Molar mass of CuSO4,
MM = 1*MM(Cu) + 1*MM(S) + 4*MM(O)
= 1*63.55 + 1*32.07 + 4*16.0
= 159.62 g/mol

use:
mass of CuSO4,
m = number of mol * molar mass
= 0.1687 mol * 1.596*10^2 g/mol
= 26.94 g
Answer: 26.9 g